If the length of the diagonal of a square is 5yes , gone the length of each side of the square

Answer by chndnbl(31)

all sides of square is equal. Draw a square and let its side be AB,BC,CD and AD and diagonal be AC.
Let AB=BC=CD=AD=a cm
Use the pythagoras theorm,
AC^2 = AB^2 + BC^2
5^2 = a^2 + a^2
25=2 a^2
a^2 = 25/2
find a=?
Check the question.